A ball is thrown down a building with an initial velocity of -5 m/s. It takes 6 seconds to reach the ground. How high is the building? What is the velocity when it hits the ground?


1
Expert's answer
2021-02-06T10:07:06-0500

Initial velocity of the ball "u=5m\/s"

Time taken by ball, "t=6s"

Final velocity is given by, "v=u+gt"

"v=5+(10*6)=65m\/s"

It is the velocity with which ball will reach the ground